OpenStudy (anonymous):

At an amusement park, each premium ride costs $12 and each gold star ride is $8. Stan spends $96 on tickets and he gets 3 more premium rides than gold star rides. How much does he spend on gold star rides? A. $3 B. $8 C. $16 D. $24

OpenStudy (dumbcow):

12P + 8G = 96 P = G+3

OpenStudy (anonymous):

Consider x = no. of premium tickets & y = no of gold star. the two simultaneous equations are: 12x + 8y = 96 x - y = 3 Solving them you get: x=6, y=3 so the answer is 3*8 = 24. The amount he spent on gold star rides.